A vendor bought toffees at 6 for a rupee. How many for a rupee must he sell to gain 20%?
A. 3
B. 4
C. 5
D. 6

Grade:8

3 Answers

Aman
5 years ago
1
Deepak Kumar Shringi
5 years ago
6 toffee = 1RS
1 toffee= 1/6 Rs
to gain 20 % 1 should sell in = 1.2 * (1/6) Rs
SP of toffee = .2 RS
in 1 rupees there will be 5 toffee he must sell

Aman
5 years ago
6 toffee =1rupee
1 toffee =1/6 rupee
Gain=20%
So sp=(1/6)×120/100 =1/5
So no of tofees in 1 rupee 1/(1/5)=5 
Hence answer =5 (c)